### v2

# Neuronal correlations shape the scaling behavior of memory capacity and nonlinear computational capability of recurrent neural networks

Abstract
Reservoir computing is a powerful framework for real-time information processing, characterized by its high computational ability and quick learning, with applications ranging from machine learning to biological systems. In this paper, we demonstrate that the memory capacity of a reservoir recurrent neural network scales sublinearly with the number of readout neurons. To elucidate this phenomenon, we develop a theoretical framework for analytically deriving memory capacity, attributing the decaying growth of memory capacity to neuronal correlations. In addition, numerical simulations reveal that once memory capacity becomes sublinear, increasing the number of readout neurons successively enables nonlinear processing at progressively higher polynomial orders. Furthermore, our theoretical framework suggests that neuronal correlations govern not only memory capacity but also the sequential growth of nonlinear computational capabilities. Our findings establish a foundation for designing scalable and cost-effective reservoir computing, providing novel insights into the interplay among neuronal correlations, linear memory, and nonlinear processing.
I Introduction
Over the past decade, deep learning has advanced significantly, with generative AI in particular beginning to transform many aspects of daily life. Although deep neural networks possess the capacity to represent highly complex functions, they typically require substantial computational resources, large volumes of training data, and long training times. These high computational demands pose a major obstacle to applying deep learning in real-time computing on edge-computing devices [1].
Reservoir computing (RC) is a machine learning framework for efficiently training large scale recurrent neural networks (RNNs), referred to as ”reservoirs” in the context of RC [2, 3]. In contrast to conventional training methods such as backpropagation through time (BPTT), RC optimizes only readout weights and leaves the remaining weights fixed, which enables quick and low-cost learning. In spite of its training simplicity, RC has high computational performance and a broad range of applications for processing time-series data [4]. RC is not restricted to RNNs, and a wide variety of dynamical systems can be utilized as reservoirs under appropriate conditions. In particular, RC using a real physical system such as soft matter and optical systems, so called ”physical reservoir computing”, has been intensively studied in recent years [5].
In a typical RC setting, the readout connections linking reservoir units to output units are sparse [5]. This means that the size of the reservoir ($N$) is substantially larger than the number of readout units ($L$) that connect to the output units. Empirically, the performance of RC improves as an increase in $L$. However, implementing a large number of readout connections can be resource-intensive, especially in physical RC implementations. Consequently, understanding the relationship between $L$ and the computational ability of a reservoir is essential for constructing cost-effective RC. In particular, performance in RC depends on both the reservoir’s memory capacity [6] and nonlinear processing capability [7], and the number of readout units must be carefully chosen to balance these two aspects according to the demands of a given task. Despite its practical importance, a practical principle for determining an appropriate $L$ remains largely unexplored. In this study, we aim to address this gap by elucidating how memory capacity and nonlinear computational ability scale with $L$, with the particular focus on the role of neuronal correlations.
We demonstrate, for the first time, that the memory capacity of an RNN increases monotonically with $L$, although its growth rate progressively declines (Fig.1). This sublinear scaling of memory capacity cannot be explained by the previous theoretical works assuming a scaling regime where $L$ is of order one relative to $N$, i.e. $L\sim O(1)$ [8, 9], because the memory capacity in this regime is proportional to $L$, as will be shown later. To bridge the gap between the observation and theory, we develop a novel theory for analytically deriving memory capacity for $L\sim O(\sqrt{N})$. Using our theoretical framework, we demonstrate that the correlations between reservoir neurons, despite their very weak magnitude scaling as $O(1/\sqrt{N})$, play an important role in the declining growth rate of memory capacity. Furthermore, numerical simulations show that the ability to perform more complex nonlinear tasks emerges supralinearly and sequentially as $L$ increases. These findings indicate that the number of readout connections should be tailored to the specific memory and nonlinearity requirements of the task at hand.
II Model setup
We investigate computational capacity of RC using a large random RNN as a reservoir (Fig.1(a)), which is known as Echo state network (ESN) [2], one of the canonical model for RC  [10, 11]. The state of $i$-th neuron at discrete time $t$ is described by the variable, $x_{i}(t)$. The activation function $\phi$ is assumed to be an odd saturating sigmoid function satisfying $\phi^{\prime}(0)=1$, $\phi^{\prime}(x)>0$, $\phi^{\prime\prime}(x)\leq 0\ (x\geq 0)$ and $\phi(\pm\infty)=\pm 1$. The time evolution of a reservoir RNN is determined by the difference equation,
$x_{i}(t)=\sum_{j=1}^{N}J_{ij}\phi(x_{j}(t-1))+u_{i}s(t)+\xi_{i}(t),$ (1)
where $N\gg 1$ indicates the total number of neurons. The recurrent weights, $J_{ij}$, and the input weights, $u_{i}$, are sampled i.i.d from Gaussian distributions, i.e. $J_{ij}\sim\mathcal{N}(0,g^{2}/N)$ and $u_{i}\sim\mathcal{N}(0,1)$. The parameter $g$ controls the strength of the recurrent weights. The input signal at time $t$, represented by $s(t)$, is Gaussian white noise with zero mean and variance $\sigma_{s}^{2}$. To take into account inherent neuronal noise, each neuron is subjected to independent Gaussian white noise, $\xi_{i}(t)$, with zero mean and variance $\sigma_{n}^{2}$. Notably, the input signal – shared across all neurons – induces correlations among reservoir neurons, whereas the independent neuronal noise acts to diminish them.
The macroscopic dynamical behavior of the large random RNN described by Eq.(1) has been extensively studied by means of dynamical mean-field theory (DMFT) [12]. In the absence of the inputs and noise, i.e. $\sigma_{s}^{2}=\sigma_{n}^{2}=0$, the RNN exhibits phase transition from zero fixed-point to chaotic dynamics at $g=1$ in the large network size limit [12, 13]. In the presence of inputs ($\sigma_{s}^{2}>0$ or $\sigma_{n}^{2}>0$), chaotic variability is typically suppressed [8, 13, 14, 15, 9, 16], although it has recently been shown that in RNNs incorporating gating mechanisms similar to those in gated recurrent units (GRUs), external inputs can instead induce chaos [17].
The output of the reservoir RNN is defined as $z(t)=\sum_{i=1}^{L}w_{i}x_{i}(t)$, where $L$ is the number of readout neurons. We assume that the readout neurons are sparse, i.e. $L\ll N$. According to the RC framework, the output weights, $\{w_{i}\}_{i=1}^{L}$, are optimized to minimize the time-averaged squared error between the output signals, $\{z(t)\}_{t=1}^{T}$, and the desired signals, $\{f(t)\}_{t=1}^{T}$.
III Memory capacity
Memory capacity [6, 7] is a commonly used benchmark for RC, quantifying how accurately a reservoir can reproduce its past Gaussian white noise inputs. Specifically, it is defined as $MC\equiv\sum_{d=0}^{\infty}M_{d}$, where
$M_{d}\equiv 1-\frac{{\rm{min}}_{\bm{w}}\langle(z(t)-s(t-d))^{2}\rangle}{\langle s(t)^{2}\rangle},$ (2)
and the angular bracket denotes time averaging. A high value of $M_{d}$ indicates that the reservoir can accurately output its input signal from $d$ steps prior. It is shown that $0\leq M_{d}\leq 1$ holds true [6, 7]. Memory capacity is defined as the infinite sum of $M_{d}$, but it has been proven to be finite and bounded by the number of readout neurons, i.e. $0\leq MC\leq L$  [6, 7]. In particular, for RC with a linear activation function operating in noise-free conditions ($\sigma_{n}^{2}=0$), memory capacity is equivalent to $L$ independently of $g$ and $\sigma_{s}^{2}$  [6].
Fig.1(b) illustrates the relationship between $L$ and the memory capacity for the reservoir RNN described by Eq.(1), obtained through numerical simulations. As shown, memory capacity exhibits monotonic growth, but its growth rate gradually diminishes as $L$ increases. Consequently, memory capacity can be characterized as a sublinear function of $L$.
Figure 1: (a) Overview of RC utilizing a random RNN receiving input signals and inherent neuronal noise. The shaded region represents the readout neurons. (b) Numerical simulations for memory capacity of the reservoir RNNs with network size $N=1000$, noise intensity $\sigma_{n}^{2}=0.1^{2}$, and activation function $\phi(x)=\tanh(x)$. Shaded area represents mean±$\pm$±std of direct numerical simulations for 10 different network and input signal realizations. The sum of $M_{d}$ is calculated up to $d=1000$. The dashed line indicates the theoretical upper bound of memory capacity, $MC=L$. Simulation time length is $T=10^{4}$.
IV Theoretical analysis of memory capacity
To elucidate the mechanism behind the decaying growth rate of memory capacity, we provide a novel theory for deriving an analytical solution of memory capacity. Our approach incorporates some concepts and methods from statistical physics. Hereafter, we make an additional assumption regarding the scale of the model parameters:
$L=\alpha\sqrt{N},\ \ \sigma_{s}^{2}=\frac{\tilde{\sigma}_{s}^{2}}{\sqrt{N}},\ \ \sigma_{n}^{2}\sim O(1),$ (3)
where both $\alpha$ and $\tilde{\sigma}_{s}^{2}$ are $O(1)$. The parameter $\alpha$ represents the relative magnitude of $L$ compared to $\sqrt{N}$. In contrast to previous studies [18, 8, 9], where $L\sim O(1)$, our model employs a significantly larger number of readout neurons, although the output connections remain sparse, as $L=\alpha\sqrt{N}\ll N$
By simple calculation, $M_{d}$, defined in Eq.(2), can be expressed as
$M_{d}=\frac{\bm{a}_{d}^{\top}C^{-1}\bm{a}_{d}}{\langle s(t)^{2}\rangle},$ (4)
where the elements of $\bm{a}_{d}\in\mathbb{R}^{L}$ and $C\in\mathbb{R}^{L\times L}$ are respectively $(\bm{a}_{d})_{i}\equiv\langle s(t-d)x_{i}(t)\rangle$ and $C_{ij}\equiv\langle x_{i}(t)x_{j}(t)\rangle$ ($i,j$ indicate the indices of the readout neurons) [7]. The inverse of the covariance matrix, $C^{-1}$, prevents us from making progress on analytical calculation. Previous theoretical studies circumvented this issue by assuming the off-diagonal entries of $C$ are zero, that is, ignoring neuronal correlations [18, 8, 9]. However, this approximation results in linear scaling of memory capacity, $MC\propto L$, conflicting with our numerical simulations as shown in Fig.1(b).
We noticed that from the scale assumption Eq.(3), the diagonal entries of $C$ are $O(1)$, whereas the non-diagonal ones are $O(1/\sqrt{N})$ (See Appendix.A.2). Hence, the non-diagonal elements are much smaller than the diagonal ones, which motivates us to perform Neumann series expansion,
$C^{-1}=(D+\tilde{C})^{-1}=\sum_{n=0}^{\infty}D^{-1}\left(-\tilde{C}D^{-1}\right)^{n},$ (5)
where $D_{ij}\equiv\delta_{ij}\langle x_{i}(t)^{2}\rangle$ and $\tilde{C}_{ij}\equiv(1-\delta_{ij})\langle x_{i}(t)x_{j}(t)\rangle$. This approach allows us to circumvent the issue of the inverse matrix and enables an analytical calculation of $M_{d}$.
In the large network size limit, we can assume self-averaging for $M_{d}$, $\lim_{N\to\infty}M_{d}=\lim_{N\to\infty}[M_{d}]$, where the square bracket denotes the average over network realizations, known as quenched average [19]. Consequently, substituting Eq.(5) into Eq.(4) and taking quenched average of $M_{d}$, we analytically obtain memory capacity:
$\lim_{N\to\infty}MC(L=\alpha\sqrt{N})=\lim_{N\to\infty}\sum_{d=0}^{\infty}\left[M_{d}(L=\alpha\sqrt{N})\right]$ (6)
$=\sum_{d=0}^{\infty}\sum_{n=0}^{\infty}(-1)^{n}\left\{\frac{\alpha\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}\left(g\langle\phi^{\prime}(x)\rangle_{x\sim\mathcal{N}(0,[\langle x_{i}^{2}\rangle])}\right)^{2d}\right\}^{n+1}$
where the value of $[\langle x_{i}^{2}\rangle]$ can be obtained by solving a self-consistent equation,
$[\langle x_{i}^{2}\rangle]=\sigma_{n}^{2}+g^{2}\langle\phi(x)^{2}\rangle_{x\sim\mathcal{N}(0,[\langle x_{i}^{2}\rangle])}.$ (7)
In deriving this result, it is necessary to calculate quenched averages of quantities involving neuronal correlations, such as
$\left[\langle s(t-d)x_{i}(t)\rangle\langle x_{i}(t)x_{j}(t)\rangle\langle s(t-d)x_{j}(t)\rangle\right]\ \ (i\neq j).$ (8)
However, these quantities cannot be evaluated using DMFT because it reduces the original $N$-body system to an effective single-body description, thereby neglecting inter-neuronal correlations. To overcome this limitation, we adopt the dynamical cavity method, a technique originally introduced by Clark et.al [20] from statistical physics to the analysis of neural networks. A detailed description of this theory is provided in Appendix.A.1 and A.2, followed by the full derivation of Eq.(6) in Appendix.B.1. In the following, we use the shorthand notation, $\langle f(x)\rangle_{*}\equiv\langle f(x)\rangle_{x\sim\mathcal{N}(0,[\langle x_{i}^{2}\rangle])}$.
[…]
$MC=\sum_{d=0}^{\infty}\sum_{n=0}^{\infty}(-1)^{n}\left\{\frac{\alpha\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}\left(\frac{g^{2}}{1+\frac{\pi}{2}[\langle x_{i}^{2}\rangle]}\right)^{d}\right\}^{n+1},$
where $[\langle x_{i}^{2}\rangle]$ is determined by solving
$[\langle x_{i}^{2}\rangle]=\sigma_{n}^{2}+g^{2}\left(-1+\frac{4}{\pi}\arctan\sqrt{1+\pi[\langle x_{i}^{2}\rangle]}\right).$
We have confirmed that the analytical values of $MC$ are consistent with the numerical simulations, exhibiting a sublinear function of $L$, for $L$ below a threshold (Fig.2). When $L$ exceeds the threshold, the analytical values rapidly deviate from the simulated ones and eventually diverge (dashed lines in Fig.2). This is because Neumann series expansion, Eq.(5), does not converge as the spectrum norm $\|\tilde{C}D^{-1}\|$ exceeds one for large $L$. To evaluate the upper bound of $L$ ensuring convergence of the series expansion, it is enough to evaluate $\|\tilde{C}D^{-1}\|$, but this is a challenging task. Instead, we derive the sufficient condition for the expansion series to converge (see Appendix.B.1), and the corresponding upper bounds of $L$ satisfying this condition are indicated by star marks in Fig.2.
[…]
V Neuronal correlations shape the sublinear scaling of memory capacity
Having obtained the analytical solution of memory capacity, we now investigate the underlying reason why memory capacity grows sublinearly with respect to $L$. To quantify the extent to which the growth rate of memory capacity decays, we define a decay rate,
$r(L)\equiv\frac{MC(L)}{L\times MC(1)}.$ (11)
By this definition, if $MC(L)$ scales linearly with $L$, then $r(L)=1$ holds for all $L$. In contrast, $r(L)$ decays from unity when $MC(L)$ is a sublinear function of $L$. As depicted in Fig.2, $r(L)$ decays for various parameter combinations, confirming the sublinear scaling of memory capacity with respect to $L$.
Employing the solution of memory capacity (Eq.(6)), we obtain an analytical form of the decay rate (see Appendix.B.2 for derivation):
$\lim_{N\to\infty}r(L=\alpha\sqrt{N})$ $=1-\sum_{n=1}^{\infty}(-1)^{n-1}\left(\frac{\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}\alpha\right)^{n}$
×$\hskip 60.0pt\times\frac{1-(g\langle\phi^{\prime}(x)\rangle_{*})^{2}}{1-(g\langle\phi^{\prime}(x)\rangle_{*})^{2n+2}}.$
It is noteworthy that the $n$-th term corresponds exactly to the $n$-th term in the Neumann series expansion, Eq.(5). Therefore, under the assumption of vanishing neuronal correlations ($\tilde{C}=O$), the summation terms in Eq.(V) vanish, resulting in $r(L)=1$. This implies that higher neuronal correlations contribute to the faster decay of $r(L)$ through the contributions of the summation terms in Eq.(V). In fact, we have confirmed that $r(L)$ decays more rapidly for hyperparameters that lead to higher neuronal correlation, such as larger input signals ($\tilde{\sigma}_{s}^{2}$), smaller recurrent weights ($g$), and lower noise intensity ($\sigma_{n}^{2}$), as depicted in Fig.3. Notably, while the slope of the memory capacity curve at the origin, which is proportional to $MC(1)$, reaches its maximum when the reservoir RNN operates near the edge of chaos [8, 9], the extent of the decay of $r(L)$ varies monotonically with the hyperparameters $g$, $\sigma_{s}^{2}$ and $\sigma_{n}^{2}$, irrespective of whether the RNN is near the edge of chaos. These results suggest that instead of proximity to the edge of chaos, neuronal correlation is a novel mechanism determining the computational ability of RC.
[…]
Thus far, we have derived the memory capacity under the specific scaling assumptions (Eq.(3)). However, practical applications often deviate from these conditions, raising the question of whether neuronal correlations remain a key determinant of sublinear memory capacity growth in more general settings. To address this, Fig.4 explores a broader range of model parameters including larger input intensities and various noise levels, and different values of $g$. Here, we introduced a half-life of memory capacity growth, $L_{\rm half}$, defined as the value of $L$ at which $r(L)$ equals $0.5$ (Fig.4, inset). A lower $L_{\rm half}$ indicates more rapid decay of $r(L)$. Notably, as shown in Fig.4, strong neuronal correlations are closely associated with a faster decay of memory capacity growth, highlighting their persistent influence even when the original scaling assumptions are relaxed.
Figure 4: Relation between the level of neuronal correlations and the half-life of memory capacity growth. Each data point corresponds to a set of hyperparameters sampled from $g\sim U(0.2,2)$, $\sigma_{s}\sim U(0.1,3)$ and $\sigma_{n}\sim U(0,3)$ (100 samples are plotted). Red points indicate parameter regimes beyond our theoretical framework, as determined by $L_{\rm half}$ surpassing the deviation threshold of the theoretical values of $MC$ (Fig.2). For each parameter set, the half-life of memory capacity growth, $L_{\rm{half}}$, and pair-wise averaged absolute values of Pearson correlation coefficients of neuronal activity, $\rho_{ij}$, are calculated across 10 network and input realizations. The mean values obtained from these 10 realizations are displayed as a single plot in the figure. The activation function is $\phi(x)=\tanh(x)$. The network size is $N=1000$, and the simulation time length is $T=10^{4}$.
VI Scaling behavior of nonlinear computational capability
While our previous analysis has focused on memory capacity, non-linear computational ability is equally crucial for RC to address complex tasks. As established in previous works, there exists a trade-off between memory capacity and nonlinear computational ability for RC  [7, 21]. Thus, we expect that nonlinear computational ability increases instead of a deceleration in the growth rate of memory capacity.
Information processing capacity (IPC) introduced by Dambre et al. [7] offers a task-independent metric for evaluating the performance of RC, enabling a comprehensive assessment of both linear and non-linear computational capabilities. Since IPC theory is somewhat complicated, we provide detailed explanations in Appendix.C.1. Put simply, the IPC for degree $D$, denoted by $IPC_{D}$, represents the reservoir’s ability to approximate $D$-th order polynomial functions of past white noise input signals. The IPC has two key properties. First, by definition, memory capacity is exactly equivalent to $IPC_{1}$, and thus, $IPC_{D}$ for $D\geq 2$ represents non-linear computational ability. Second, it has been proven [7] that the total IPC, defined as $IPC_{\rm total}\equiv\sum_{D=1}^{\infty}IPC_{D}$, equals to the number of readout units, $L$, provided that the reservoir satisfies the echo state property, a necessary condition for a reservoir to be usable for RC. Here, the echo state property implies that the reservoir’s state is uniquely determined solely by input signals regardless of reservoir’s initial state [2].
Fig.5 illustrates how the values of $IPC_{D}$ for our RNN model (Eq.(1)) change with increasing $L$ across various hyperparameter combinations ($g$, $\sigma_{s}^{2}$, and $\sigma_{n}^{2}$), as obtained through numerical simulations. Note that, due to the symmetry of the model, $IPC_{D}$ values for even-degree $D$ are identically zero. As shown, a decline in the growth of the $IPC_{1}$ (equivalent to memory capacity) is accompanied by an increase in the non-linear computational ability represented by $IPC_{D}$ for $D\geq 2$, reflecting memory-nonlinearity trade-off.
Our novel finding is that, as $L$ increases, successively higher-degree IPCs are sequentially activated from lower to higher orders, accompanied by supralinear rises in the reservoir’s nonlinear computational capabilities. This phenomenon is observed not only in an RNN satisfying the echo state property (Fig.5(a)), but also in a chaotic RNN (Fig.5(b)) and in an RNN subject to neuronal noise (Fig.5(c)), where the echo state property breaks down. Additionally, the same finding is observed for an RNN employing the ReLU activation function (Fig.8), where our theoretical framework for deriving memory capacity is not applicable (see Appendix.C.2).
As analytical derivation of higher-degree IPC is beyond our capability, the detailed mechanism underlying these observations remains unclear. However, it is plausible that the neuronal correlation plays a pivotal role, analogous to its influence on memory capacity. This hypothesis is supported by speculation that each $IPC_{D}$ would exhibit linear scaling with $L$ if the effects of neuronal correlations were neglected, which contrasts with the observed supralinear and sequential emergence of $IPC_{D}$ (see Appendix.C.3).
Figure 5: IPC values for varying the number of readout neurons $L$ of a reservoir RNN with activation function $\phi(x)=\tanh(x)$ and network size $N=1000$. The top row panels show stacked representations of IPC values for each degree ($IPC_{D}$), while the bottom row panels provides a detailed breakdown of individual $IPC_{D}$ values. An enlarged view of the smaller IPC values is shown in the insets in the bottom panels. The values of IPC for up to degree 9 are obtained. The model parameters $(g,\sigma_{s},\sigma_{n})$ are $(0.9,0.3,0.0)$ for (a), $(2.5,2.0,0.0)$ for (b), and $(0.8,0.3,0.1)$ for (c), respectively. While $\sum_{D=1}^{9}IPC_{D}$ almost equals $L$ for (a), it falls significantly short of $L$ for (b) and (c) (a dashed line denotes $\sum_{D=1}^{\infty}IPC_{D}=L$). This is attributed to the violation of the echo state property. Specifically, the RNN for (b) is chaotic, as indicated by a positive maximum conditional Lyapunov exponent [22, 15] ($\lambda_{\rm max}=0.087\pm 0.0067$), while the RNN for (c) is subject to noise. Shaded area represents mean±$\pm$±std for 10 different network and input signal realizations. Simulation time length is $T=10^{5}$.
VII Discussion
In the present study, we have investigated the relationship between the number of readout neurons and the computational capacity for large random RNN reservoirs. Through analytical and numerical approaches, we have demonstrated that the memory capacity grows sublinearly with $L$, and that nonlinear computational capabilities emerge incrementally with $L$, enabling the reservoir to handle increasingly complex tasks. The specific memory and nonlinearity requirements of a given task can be assessed using the framework proposed by [23]. In this context, our results provide a practical principle for designing cost-effective RC systems: while increasing $L$ raises computational and structural costs and yields diminishing returns for memory, it becomes essential when tasks demand higher-order nonlinear computations.
Our findings have implications that extend beyond the framework of RC. First, in neuroscience, the impacts of neuronal correlations on function have been extensively investigated from the perspective of neural manifold [24, 25, 26]. A neuronal network with a high level of correlations occupies a restricted region of $N$-dimensional phase space, whose linear dimension is quantified by effective dimension [27]. In the context of population coding, a well-established trade-off exists between representation richness and robustness: high-dimensional (low-correlation) activity enables complex and diverse information encoding, while low-dimensional (high-correlation) activity supports redundant encoding and noise robustness [25, 26]. However, the mammalian cortex in the awake state is known to operate in an asynchronous regime characterized by near-zero neuronal correlations [28, 29, 30], akin to those observed in our large random RNNs. In addition, these prior studies have only focused on the implications of neuronal correlations for static information processing, i.e., mapping static inputs to outputs. Our findings suggest a novel role for weak correlations in dynamic information processing, offering a new hypothesis about their functional significance in the brain.
Second, in the field of deep learning theory, it has been established that neural networks trained via gradient-based methods, such as BPTT, can learn tasks both in the lazy regime, where their weights undergo only infinitesimal changes from their initial random values [31, 32], and in the rich regime, where weights change substantially to extract task-relevant features [33, 34, 35]. Our insights may provide valuable understanding of the computational capacity of RNNs trained in the lazy regime, as their recurrent weights remain nearly identical to those of untrained, reservoir-style networks.
In conclusion, our study not only advances the understanding of RC but also sheds light on the broader implications of neuronal correlations and their effects on both biological and artificial systems. These insights offer new directions for optimizing RNN architectures and exploring the fundamental nature of computation in complex networks.
Acknowledgements.
We thank Joni Dambre for providing the source code for computing information processing capacity, and Tomoyuki Kubota for helpful discussions regarding its implementation. S.T. was supported by JSPS KAKENHI Grant No. JP22KJ1959. T.A. was supported by MEXT KAKENHI Grant Numbers 23H04467 and by JSPS KAKENHI Grant Numbers 24H00723, 20K20520.
Appendix A THEORETICAL ANALYSIS ON STATISTICS OF NEURONAL ACTIVITY
In this section, we derive statistics of neuronal activity based on dynamical cavity approach [20, 36]. Several results and methodologies introduced here are frequently utilized in deriving memory capacity in Appendix.B.
In deriving single neuronal statistics such as the variance of $x_{i}(t)$ (see Appendix.A.1) in the limit of $N\to\infty$, the same result can be obtained via dynamical mean-field theory [12, 19] . However, since the dynamical mean-field theory transforms the $N$-body system into the effective single-body system, it is inherently unable to evaluate many-body neuronal statistics, such as neuronal correlations (see Appendix.A.2) and memory capacity(see Appendix.B.2). In contrast, the dynamical cavity method is, in principle, applicable to $n$-body neuronal statistics for any $n$. Furthermore, while the aforementioned limitation of dynamical mean-field theory has recently been addressed through subleading corrections to the saddle-point solution [37, 38], the dynamical cavity method remains advantageous for its conceptual clarity.
In the following, we use the shorthand notation for brevity $x_{i,t}\equiv x_{i}(t)$, $s_{t}\equiv s(t)$, $\phi_{i,t}\equiv\phi(x_{i,t})$ and $\langle x_{i}x_{j}\rangle\equiv\langle x_{i}(t)x_{j}(t)\rangle$.
Before delving into the details, we first outline the necessity of employing the dynamical cavity method. Consider calculating the quenched variance of neuronal activity, $[x_{i,t}^{2}]$. For simplicity, we ignore input and noise signals. According to time evolution equation of neurons (Eq.(1)), $x_{i,t}$ is the sum of a large number ($N\gg 1$) of mutually independent quantities, $J_{ij}\phi(x_{j,t-1})$. This implies that $x_{i,t}$ follows a Gaussian distribution due to the central limit theorem, and thus, it is sufficient to determine the complete statistics of $x_{i,t}$ by calculating its mean and variance. Given the symmetry of the model, we have $[x_{i,t}^{2}]=0$. On the other hand, we can intuitively calculate $[x_{i,t}^{2}]$ as:
$[x_{i,t}^{2}]=\sum_{j,k}^{N}\left[J_{ij}J_{ik}\phi_{j,t-1}\phi_{k,t-1}\right]\overset{?}{=}\sum_{j,k}^{N}\left[J_{ij}J_{ik}\right]\left[\phi_{j,t-1}\phi_{k,t-1}\right]=g^{2}\left[\phi_{j,t-1}^{2}\right]=g^{2}\langle\phi(x)\rangle_{x\sim\mathcal{N}(0,[x_{i,t-1}^{2}])},$ (13)
yielding time series of $[x_{i,t}^{2}]$. Here we assumed in the second equation that the recurrent weights, $J_{ij}$ and $J_{ik}$, are independent of the neuronal activities, $\phi_{j}$ and $\phi_{k}$. Fortunately, this assumption is validated in the limit of $N\to\infty$, as shown with the dynamical mean-field theory. However, this assumption is likely to break down when finite size effects are taken into account. To circumvent this challenge, we need to employ the dynamical cavity method.
Figure 6: Overview of dynamical cavity method. For simplicity, only one auxiliary neuron is illustrated, but in practice, $n$ auxiliary neurons are introduced to calculate $n$-body statistics. Latin indices, $i,j,\cdots$, are used to label reservoir neurons, while Greek indices, $\mu,\nu,\cdots$, denote auxiliary neurons. The weights connecting auxiliary neurons to reservoir neurons, as well as the self-connections of auxiliary neurons, are sampled from the same distribution as the recurrent weights of the reservoir.
A.1 Single neuronal statistics
We aim to derive the quenched average of the time-variance of neuronal activity, $[\langle x_{i}^{2}\rangle]$. Note that due to the self-averaging, it is equivalent to the population average of time-variance to leading order, i.e. $[\langle x_{i}^{2}\rangle]=\frac{1}{N}\sum_{i}^{N}\langle x_{i}^{2}\rangle$. The symmetry of the model setting ensures that the quenched average of the time-mean, $[\langle x_{i}\rangle]$, vanishes.
The overview of dynamical cavity method is described in Fig.6. One auxiliary neuron, indexed by 00, is added to the original reservoir RNN whose neurons are indexed by $i=1,\cdots,N$. The original neuron is perturbated by the addition of the auxiliary neuron. By denoting the perturbation applied to the $i$th neuron at time $t$ by $\delta_{i,t}$, we describe the dynamics of the auxiliary neuron as
$x_{0,t+1}$ $=\sum_{i=1}^{N}J_{0i}\phi(x_{i,t}+\delta_{i,t})+J_{00}\phi_{0,t}+u_{0}s_{t+1}+\xi_{0,t+1}$ (14)
$=\sum_{i=1}^{N}J_{0i}\phi_{i,t}+\sum_{i=1}^{N}J_{0i}\phi^{\prime}_{i,t}\delta_{i,t}+J_{00}\phi_{0,t}+u_{0}s_{t+1}+\xi_{0,t+1}.$
Here, it should be noted that $x_{i,t}$ represents the original activity of a reservoir neuron before the introduction of the auxiliary neuron. Therefore, the value of $J_{0i}$ is completely independent of reservoir activities, $x_{i,t}$ and $\phi_{i,t}$, in contrast to the intuitive argument in Eq.(13).
We define $\chi_{ij,ts}\equiv\delta x_{i,t}/\delta I_{j,s}$ as the linear susceptibility of the $i$th neuron at time t, induced by an infinitesimal external inputs applied to the $j$th neuron at time $s\leq t$, $I_{j,s}$. Then, as the perturbation to $i$th neuron at time $t$, $\delta_{i,t}$, is the sum of all direct perturbations to $j\in\{1,\cdots,N\}$ at time $s<t$ weighted by the corresponding susceptibility $\chi_{ij,ts}$, we obtain
$\delta_{i,t}=\sum_{s<t}\sum_{j=1}^{N}\chi_{ij,ts}J_{j0}\phi_{0,s}.$ (15)
We substitute Eq.(15) into Eq.(14), yielding
$x_{0,t+1}$ $=\underbrace{\sum_{i=1}^{N}J_{0i}\phi_{i,t}}_{\equiv\eta_{t}}+\sum_{s<t}\underbrace{\sum_{i,j}^{N}\chi_{ij,ts}J_{0i}J_{j0}\phi^{\prime}_{i,t}}_{\equiv\kappa_{s}}\phi_{0,s}+J_{00}\phi_{0,t}+u_{0}s_{t+1}+\xi_{0,t+1}$
$\equiv\eta_{t}+\sum_{s<t}\kappa_{s}\phi_{0,s}+J_{00}\phi_{0,t}+u_{0}s_{t+1}+\xi_{0,t+1}.$
The physical interpretation of this equation is as follows. The first term, $\eta_{t}$, is the feedforward input from the reservoir. The second term, $\kappa_{s}\phi_{0,s}$, arises from the perturbation induced in the reservoir by the auxiliary neuron 00, which is read out by neuron 00. The third term originates from self-connection of the auxiliary neuron 00.
The order of the first term is evaluated as
$\left[\eta_{t}^{2}\right]=\sum_{i,j}^{N}[J_{0i}J_{0j}\phi_{i,t}\phi_{j,t}]=\sum_{i,j}^{N}[J_{0i}J_{0j}][\phi_{i,t}\phi_{j,t}]=g^{2}[\phi^{2}_{i,t}]\sim O(1)$ (17)
Here we used in the second equation the fact that $J_{0i}$ and $J_{0j}$ are independent of $\phi_{i,t}$ and $\phi_{j,t}$ as mentioned above. The order of the third term is evaluated as
$\left[\left(J_{00}\phi_{0,t}\right)^{2}\right]\leq[J_{00}^{2}]\sim O(1/N).$ (18)
For evaluation of the second term in Eq.(A.1), we assess the order of $\chi_{ij,ts}$. We begin with $s=t-1$. Since $\chi_{ij,t\hskip 1.0ptt-1}=J_{ij}\phi^{\prime}_{j,t-1}$, we find $[\chi_{ij,t\hskip 1.0ptt-1}^{2}]\leq[J_{ij}^{2}]\sim O(1/N)$. Subsequently, for $s=t-2$, $\chi_{ij,t\hskip 1.0ptt-2}=\sum_{k}^{N}J_{ik}\phi^{\prime}_{k,t-1}\chi_{kj,t-1\hskip 1.0ptt-2}$, leading to $[\chi_{ij,t\hskip 1.0ptt-2}^{2}]\leq\sum_{k}^{N}[J_{ik}^{2}][(\chi_{kj,t-1\hskip 1.0ptt-2})^{2}]\sim O(1/N)$, where we used the result for $s=t-1$, $[(\chi_{kj,t-1\hskip 1.0ptt-2})^{2}]\sim O(1/N)$. By inductively repeating this analysis, we conclude that $[\chi_{ij,ts}^{2}]\sim O(1/N)$ for any $s<t$. Therefore, the order of the second term is evaluated as
$[\kappa_{s}^{2}]$ $=\sum_{i,j,k,l}^{N}[J_{0i}J_{j0}J_{0k}J_{l0}][\phi^{\prime}_{i,t}\phi^{\prime}_{k,t}\chi_{ij,ts}\chi_{kl,ts}]\leq\frac{g^{4}}{N^{2}}\sum_{i,j}^{N}[\chi_{ij,ts}^{2}]\sim O(1/N).$ (19)
Eventually, we obtain $[\kappa_{s}^{2}]\sim O(1/N)$ for $s<t$.
Employing Cauchy–Schwarz inequality, it is shown that all cross terms between $\eta_{t}$, $\kappa_{s}$, and $J_{00}\phi_{0,t}$ are $O(1/\sqrt{N})$.
Leveraging all these order estimates, from Eq.(A.1), we obtain
$[x_{0,t+1}^{2}]$ $=\left[\big{(}\eta_{t}+\sum_{s<t}\kappa_{s}\phi_{0,s}+J_{00}\phi_{0,t}+u_{0}s_{t+1}+\xi_{0,t+1}\big{)}^{2}\right]$ (20)
$=[\eta_{t}^{2}]+[u_{0}^{2}]s_{t+1}^{2}+[\xi_{0,t+1}^{2}]+O(1/\sqrt{N})$
$=g^{2}[\phi_{i,t}^{2}]+s_{t+1}^{2}+\sigma_{n}^{2}+O(1/\sqrt{N}).$
Thus, to leading order, we can ignore the non-trivial terms arising from the effects of the perturbations induced by the auxiliary neuron. The statistical behavior of the auxiliary neuron is equivalent to that of the reservoir neurons, which enables the replacement of $[x_{0,t}^{2}]$ with $[x_{i,t}^{2}]$, resulting in
$[x_{i,t+1}^{2}]=g^{2}[\phi_{i,t}^{2}]+s_{t+1}^{2}+\sigma_{n}^{2}+O(1/\sqrt{N}).$ (21)
For the scailing assumption given by Eq.(3), we can assume the inputs to be subleading, obtaining
$[x_{i,t+1}^{2}]=g^{2}[\phi_{i,t}^{2}]+\sigma_{n}^{2}+O(1/\sqrt{N}).$ (22)
Taking the time average and the network size limit ($N\to\infty$) yields a dynamical mean-field equation,
$[\langle x_{i}^{2}\rangle]=g^{2}[\langle\phi_{i}^{2}\rangle]+\sigma_{n}^{2},$ (23)
which is identical to Eq.(7) in the main text.
For later use, we evaluate the order of the quenched variance of $\langle x_{i}^{2}\rangle$, denoted by ${\rm{Var}}[\langle x_{i}^{2}\rangle]$, which is equivalent to the population variance because of the self-averaging property. Since the cavity method allows us to ignore the effects of perturbation induced by the auxiliary neuron, we easily calculate $[\langle x_{0}^{2}\rangle^{2}]$ as
$[\langle x_{0}^{2}\rangle^{2}]$ $=\left[\left\langle\left(\sum_{i}^{N}J_{0i}\phi_{i,t-1}+u_{0}s_{t}+\xi_{0,t}\right)^{2}\right\rangle^{2}\right]$ (24)
$=\frac{2g^{4}}{N^{2}}\sum_{i,j}^{N}[\langle\phi_{i}\phi_{j}\rangle^{2}]+\frac{g^{4}}{N^{2}}\sum_{i,j}^{N}[\langle\phi_{i}^{2}\rangle\langle\phi_{j}^{2}\rangle]+2g^{2}(\sigma_{s}^{2}+\sigma_{n}^{2})[\langle\phi_{i}^{2}\rangle]+3\sigma_{s}^{4}+\sigma_{n}^{4}$
$=2g^{4}[\langle\phi_{i}\phi_{j}\rangle^{2}]_{i\neq j}+\frac{3g^{4}}{N}[\langle\phi_{i}^{2}\rangle^{2}]+g^{4}[\langle\phi_{i}^{2}\rangle\langle\phi_{j}^{2}\rangle]_{i\neq j}+2g^{2}(\sigma_{s}^{2}+\sigma_{n}^{2})[\langle\phi_{i}^{2}\rangle]+3\sigma_{s}^{4}+\sigma_{n}^{4}.$
Employing Eq.(21) and Eq.(24), we obtain
${\rm{Var}}[\langle x_{0}^{2}\rangle]$ $=[\langle x_{0}^{2}\rangle^{2}]-[\langle x_{0}^{2}\rangle]^{2}$ (25)
$=g^{4}\left([\langle\phi_{i}^{2}\rangle\langle\phi_{j}^{2}\rangle]_{i\neq j}-[\langle\phi_{i}^{2}\rangle]^{2}\right)+2g^{4}[\langle\phi_{i}\phi_{j}\rangle^{2}]_{i\neq j}+\frac{3g^{4}}{N}[\langle\phi_{i}^{2}\rangle^{2}]+2\sigma_{s}^{4}.$
Specifically, when $\sigma_{s}^{2}=\tilde{\sigma}_{s}^{2}/\sqrt{N}$, we have
${\rm{Var}}[\langle x_{0}^{2}\rangle]$ $=g^{4}\left([\langle\phi_{i}^{2}\rangle\langle\phi_{j}^{2}\rangle]_{i\neq j}-[\langle\phi_{i}^{2}\rangle]^{2}\right)+2g^{4}[\langle\phi_{i}\phi_{j}\rangle^{2}]_{i\neq j}+\frac{3g^{4}}{N}[\langle\phi_{i}^{2}\rangle^{2}]+\frac{2\tilde{\sigma}_{s}^{4}}{N}.$ (26)
As shown in Appendix.A.2, we see $[\langle\phi_{i}\phi_{j}\rangle^{2}]_{i\neq j}\sim O(1/N)$. The self-averaging property allows us to evaluate the first term as
$[\langle\phi_{i}^{2}\rangle\langle\phi_{j}^{2}\rangle]_{i\neq j}-[\langle\phi_{i}^{2}\rangle]^{2}\sim\frac{1}{N^{2}-N}\sum_{i,j\ (i\neq j)}^{N}\langle\phi_{i}^{2}\rangle\langle\phi_{j}^{2}\rangle-\left(\frac{1}{N}\sum_{i}^{N}\langle\phi_{i}^{2}\rangle\right)^{2}\sim O\left(\frac{1}{N}\right).$ (27)
Consequently, we obtain ${\rm{Var}}[\langle x_{i}^{2}\rangle]\sim O(1/N)$.
A direct consequence of the scaling ${\rm{Var}}[\langle x_{i}^{2}\rangle]\sim O(1/N)$ is the following equation to leading order:
$\left[\frac{1}{\langle x_{i}^{2}\rangle}\right]=\frac{1}{[\langle x_{i}^{2}\rangle]},$ (28)
provided that $[\langle x_{i}^{2}\rangle]\sim O(1)$.
A.2 Statistics for neuronal correlations
We aim to derive the quenched average of the squared neuronal correlation, $[\langle x_{i}x_{j}\rangle^{2}]$, which is equivalent to the population average to leading order, $[\langle x_{i}x_{j}\rangle^{2}]=\frac{1}{N^{2}-N}\sum_{i,j(i\neq j)}^{N}\langle x_{i}x_{j}\rangle^{2}$. Similarly to the single neuronal statistics, the model’s symmetry ensures that $[\langle x_{i}x_{j}\rangle]=0$.
As in Appendix.A.1, we introduce two auxiliary neurons indexed by 00 and $0^{\prime}$ to the original reservoir RNN. Following the notation in Clark’s work [20], Latin indices are used to label reservoir neurons, while Greek indices denote auxiliary neurons. The dynamics of the auxiliary neurons $\mu\in\{0,0^{\prime}\}$ is described by
$x_{\mu,t+1}$ $=\sum_{i}^{N}J_{\mu i}\phi(x_{i,t}+\delta_{i,t})+\sum_{\nu}J_{\mu\nu}\phi_{\nu,t}+u_{\mu}s_{t+1}+\xi_{\mu,t+1}$ (29)
$=\sum_{i}^{N}J_{\mu i}\phi_{i,t}+\sum_{i}^{N}J_{\mu i}\phi^{\prime}_{i,t}\delta_{i,t}+\sum_{\nu}J_{\mu\nu}\phi_{\nu,t}+u_{\mu}s_{t+1}+\xi_{\mu,t+1}.$
The perturbation $\delta_{i,t}$ is given by
$\delta_{i,t}=\sum_{\nu}J_{i\nu}\phi_{\nu,t}+\sum_{s<t}\sum_{\nu}\sum_{j}^{N}\chi_{ij,ts}J_{j\nu}\phi_{\nu,s}.$ (30)
Substituting Eq.(30) into Eq.(29), we obtain
$x_{\mu,t+1}$ $=\underbrace{\sum_{i}^{N}J_{\mu i}\phi_{i,t}}_{\equiv\eta_{\mu,t}}+\sum_{s<t}\sum_{\nu}\underbrace{\sum_{i,j}^{N}J_{\mu i}J_{j\nu}\phi^{\prime}_{i,t}\chi_{ij,ts}}_{\equiv\kappa_{\mu\nu,ts}}\phi_{\nu,s}+\sum_{\nu}J_{\mu\nu}\phi_{\nu,t}+u_{\mu}s_{t+1}+\xi_{\mu,t+1}$
$\equiv\eta_{\mu,t}+\sum_{\nu}\sum_{s<t}\kappa_{\mu\nu,ts}\phi_{\nu,s}+\sum_{\nu}J_{\mu\nu}\phi_{\nu,t}+u_{\mu}s_{t+1}+\xi_{\mu,t+1}.$
The physical interpretation of this equation is analogous to the case of single neuronal statistics (Eq.A.1): $\eta_{\mu,t}$ denotes the direct feedforward input from the reservoir to the auxiliary neuron $\mu$, and $\kappa_{\mu\nu,ts}\phi_{\nu,s}$ arises from the perturbation induced in the reservoir by the auxiliary neuron $\nu$ at time $s$, which is readout by the auxiliary neuron $\mu$ at time $t$.
Expanding $[\langle x_{0}x_{0^{\prime}}\rangle^{2}]$ using Eq.(A.2), we need to evaluate all cross terms such as $[\langle\eta_{0,t}\eta_{0^{\prime},t}\rangle^{2}]$ and $[\langle\kappa_{0\mu,ts}\kappa_{0^{\prime}\nu,ts}\rangle^{2}]$. We can accomplish this task through direct calculation, analogous to the approach used in Appendix.A.1. First,
$[\langle\eta_{0,t}\eta_{0^{\prime},t}\rangle^{2}]$ $=\sum_{i,j,k,l}^{N}\left[J_{0i}J_{0^{\prime}j}J_{0k}J_{0^{\prime}l}\langle\phi_{i}\phi_{j}\rangle\langle\phi_{k}\phi_{l}\rangle\right]=\sum_{i,j,k,l}^{N}[J_{0i}J_{0^{\prime}j}J_{0k}J_{0^{\prime}l}][\langle\phi_{i}\phi_{j}\rangle\langle\phi_{k}\phi_{l}\rangle]=\frac{g^{4}}{N}[\langle\phi_{i}^{2}\rangle^{2}]+g^{4}[\langle\phi_{i}\phi_{j}\rangle^{2}]_{(i\neq j)},$
where we used in the second equation the fact that the weights connecting to auxiliary neurons are independent from the reservoir neurons’ activities. By making an ansatz, $[\langle\phi_{i}\phi_{j}\rangle^{2}]_{(i\neq j)}\sim O(1/N)$, we obtain $[\langle\eta_{0,t}\eta_{0^{\prime},t}\rangle^{2}]\sim O(1/N)$. Next, in the same manner, we obtain
$[\langle\kappa_{0\mu,ts}\kappa_{0^{\prime}\nu,ts}\rangle^{2}]$ $=\frac{g^{8}}{N^{4}}\sum_{i,j,k,l}^{N}[\langle\phi^{\prime}_{i,t}\phi^{\prime}_{k,t}\chi_{ij,ts}\chi_{kl,ts}\rangle^{2}]\leq g^{8}[\langle(\phi^{\prime}_{i,t})^{2}\rangle]^{2}[\langle\chi_{ij,ts}^{2}\rangle]^{2}\sim O(1/N^{2}).$
Repeating the same calculations, all cross terms except for the two $O(1/N)$ terms, $[\langle\eta_{0,t}\eta_{0^{\prime},t}\rangle^{2}]$ and $[\langle(u_{0}s_{t})(u_{0^{\prime}}s_{t})\rangle^{2}]=\tilde{\sigma}_{s}^{4}/N$, are shown to scale as $O(1/N^{2})$. Consequently,
$[\langle x_{0}x_{0^{\prime}}\rangle^{2}]=\dfrac{g^{4}}{N}[\langle\phi_{i}^{2}\rangle^{2}]+g^{4}[\langle\phi_{i}\phi_{j}\rangle^{2}]_{(i\neq j)}+\frac{\tilde{\sigma}_{s}^{4}}{N}+O\left(\frac{1}{N^{2}}\right).$
Replacing $[\langle x_{0}x_{0^{\prime}}\rangle^{2}]$ with $[\langle x_{i}x_{j}\rangle^{2}]_{(i\neq j)}$, we obtain
$[\langle x_{i}x_{j}\rangle^{2}]=\dfrac{g^{4}}{N}[\langle\phi_{i}^{2}\rangle^{2}]+g^{4}[\langle\phi_{i}\phi_{j}\rangle^{2}]+\frac{\tilde{\sigma}_{s}^{4}}{N}+O\left(\frac{1}{N^{2}}\right),$
for $i\neq j$. From the ansatz, $[\langle\phi_{i}\phi_{j}\rangle^{2}]_{(i\neq j)}\sim O(1/N)$, we see that $[\langle x_{i}x_{j}\rangle^{2}]_{(i\neq j)}\sim O(1/N)$.
Consequently, similar to single-neuron statistics, the non-trivial perturbation induced by auxiliary neurons, i.e., the second and third terms in Eq.(A.2), does not contribute to the leading order of neuronal correlations. In contrast, in the case of a continuous-time RNN, the perturbation induced by auxiliary neurons contributes to the leading order of neuronal correlations [20]. This discrepancy stems from the difference in the strength of the autocorrelation, $\langle x_{i}(t)x_{i}(t-\tau)\rangle_{t}$, which scales as $O(1)$ in the continuous-time model, but as $O(1/\sqrt{N})$ in our discrete-time model. This property of negligible perturbation effects by auxiliary neurons in a discrete-time model greatly simplifies the analytical derivation of memory capacity (see Appendix.B.1). Note that even in a discrete-time model, perturbation-induced effects contribute to the leading-order term of neuronal correlations if the RNN dynamics include a leak term, i.e. $x_{i}(t+1)=\gamma x_{i}(t)+\sum_{j}^{N}J_{ij}\phi_{j}(t)$ with $\gamma>0$, which yields an $O(1)$ autocorrelation.
To calculate $[\langle\phi_{i}\phi_{j}\rangle^{2}]$, we introduce Price’s theorem [19]. For two Gaussian variables, $\begin{pmatrix}z_{1}\\ z_{2}\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}c_{1}&\tau\\ \tau&c_{2}\\ \end{pmatrix}\right)$, and any smooth activation function $\phi$, we define $f_{\phi}(c_{1},c_{2},\tau)\equiv\langle\phi(z_{1})\phi(z_{2})\rangle$. Price’s theorem then states:
$\partial_{\tau}f_{\phi}(c_{1},c_{2},\tau)=f_{\phi^{\prime}}(c_{1},c_{2},\tau)$ (36)
Iteratively applying this theorem yields $\partial_{\tau}^{n}f_{\phi}(c_{1},c_{2},\tau)=f_{\phi^{(n)}}(c_{1},c_{2},\tau)$, which allows us to expand $\langle\phi(z_{1})\phi(z_{2})\rangle$ around $\tau=0$, resulting in
$\langle\phi(z_{1})\phi(z_{2})\rangle$ $=\sum_{n\geq 0}\frac{1}{n!}\tau^{n}f_{\phi^{(n)}}(c_{1},c_{2},0)=\sum_{n\geq 0}\frac{1}{n!}\langle z_{1}z_{2}\rangle^{n}\langle\phi^{(n)}(z_{1})\rangle\langle\phi^{(n)}(z_{2})\rangle.$ (37)
Applying Eq.(37) to $\langle\phi_{i}\phi_{j}\rangle$ yields
$\langle\phi_{i}\phi_{j}\rangle$ $=\langle\phi^{\prime}_{i,t}\rangle\langle\phi^{\prime}_{j,t}\rangle\langle x_{i}x_{j}\rangle+\frac{1}{3!}\langle\phi^{(3)}_{i,t}\rangle\langle\phi^{(3)}_{j,t}\rangle\langle x_{i}x_{j}\rangle^{3}+\frac{1}{5!}\langle\phi^{(5)}_{i,t}\rangle\langle\phi^{(5)}_{j,t}\rangle\langle x_{i}x_{j}\rangle^{5}+\cdots,$ (38)
where we used $\langle\phi^{(2n)}(x)\rangle=0$ since $\phi^{(2n)}$ is an odd function. Therefore, we obtain
$[\langle\phi_{i}\phi_{j}\rangle^{2}]=[\langle\phi^{\prime}_{i,t}\rangle^{2}]^{2}[\langle x_{i}x_{j}\rangle^{2}]+O\left(\frac{1}{N^{2}}\right).$ (39)
Substituting Eq.(39) into Eq.(A.2) yields
$[\langle x_{i}x_{j}\rangle^{2}]$ $=\dfrac{g^{4}}{N}\langle\phi(x)^{2}\rangle_{*}^{2}+g^{4}\langle\phi^{\prime}(x)\rangle_{*}^{4}[\langle x_{i}x_{j}\rangle^{2}]+\frac{\tilde{\sigma}_{s}^{4}}{N}+O\left(\frac{1}{N^{2}}\right),$
where $\langle f(x)\rangle_{*}$ denotes taking the average of $f(x)$ where $x\sim\mathcal{N}(0,[\langle x_{i}^{2}\rangle])$. Therefore, we obtain
$[\langle x_{i}x_{j}\rangle^{2}]=\frac{1}{N}\frac{g^{4}\langle\phi(x)^{2}\rangle_{*}^{2}+\tilde{\sigma}_{s}^{4}}{1-g^{4}\langle\phi^{\prime}(x)\rangle_{*}^{4}}+O\left(\frac{1}{N^{2}}\right),$ (41)
which is self-consistent with the ansatz, $[\langle\phi_{i}\phi_{j}\rangle^{2}]\sim O(1/N)$.
A.3 Proof of $10<g\langle\phi^{\prime}(x)\rangle_{*}<$ 1
We prove an inequality, $10<g\langle\phi^{\prime}(x)\rangle_{*}<$ 1 for any $g>0$, $\sigma_{s}^{2}\geq 0$, $\sigma_{n}^{2}\geq 0$, and an odd saturated sigmoid function satisfying $\phi^{\prime}(0)=1$, $\phi^{\prime}(x)>0$, $\phi^{\prime\prime}(x)\leq 0\ (x\geq 0)$ and $\phi(\pm\infty)=\pm 1$. We use the abbreviation, $K\equiv[\langle x_{i}^{2}\rangle]$. It is evident that $g\langle\phi^{\prime}(x)\rangle_{*}>0$ because $g>0$ and $\phi^{\prime}(x)\geq 0$. We prove that $g\langle\phi^{\prime}(x)\rangle_{*}<1$ below.
$\langle\phi^{\prime}(x)\rangle_{*}$ is a decreasing function of $K$, because
$\frac{d}{dK}\langle\phi^{\prime}(x)\rangle_{*}$ $=\frac{d}{dK}\int_{-\infty}^{\infty}\phi^{\prime}(\sqrt{K}z)Dz=\frac{1}{\sqrt{K}}\int_{0}^{\infty}z\phi^{\prime\prime}(\sqrt{K}z)Dz\leq 0,$ (42)
where $Dz$ denotes a normal Gaussian measure. Since $K$ increases with both $\sigma_{n}^{2}$ and $\sigma_{s}^{2}$, it follows that $g\langle\phi^{\prime}(x)\rangle_{*}$ decreases correspondingly. Therefore, it is enough to show that $g\langle\phi^{\prime}(x)\rangle_{*}\leq 1$ only for the case of no input and noise, $\sigma_{n}^{2}=\sigma_{s}^{2}=0$. In this case, from Eq.(7), $K$ is determined by solving
$K=g^{2}\int\phi^{2}(\sqrt{K}z)Dz.$ (43)
For $g\leq 1$, the only solution of this equation is $K=0$, resulting in $g\langle\phi^{\prime}(x)\rangle_{*}=g\phi^{\prime}(0)\leq 1$. For $g>1$, $K$ is uniquely determined as a function of $g$, so that $g$ can be assumed to be a function of $K$,
$g^{2}=\frac{K}{\int\phi^{2}(\sqrt{K}z)Dz}.$ (44)
Therefore, it suffices to show the inequality,
$g^{2}\left(\int\phi^{\prime}(\sqrt{K}z)Dz\right)^{2}=\frac{K\left(\int\phi^{\prime}(\sqrt{K}z)Dz\right)^{2}}{\int\phi^{2}(\sqrt{K}z)Dz}\leq 1,$
for any $K>0$. Employing integration by parts and Cauchy-Schwarz inequality, we can prove this inequality as
$K\left(\int\phi^{\prime}(\sqrt{K}z)Dz\right)^{2}$ $=\left(\int z\phi(\sqrt{K}z)Dz\right)^{2}\leq\int z^{2}Dz\int\phi^{2}(\sqrt{K}z)Dz=\int\phi^{2}(\sqrt{K}z)Dz.$ (46)
Appendix B ANALYSIS ON MEMORY CAPACITY
B.1 Analytical derivation of memory capacity
In this subsection, we derive the analytical solution of memory capacity. By simple calculation, $M_{d}$ defined in Eq.(2) can be expressed as
$M_{d}=\frac{\bm{a}_{d}^{\top}C^{-1}\bm{a}_{d}}{\langle s(t)^{2}\rangle},$ (47)
where the elements of $\bm{a}\in\mathbb{R}^{L}$ and $C\in\mathbb{R}^{L\times L}$ are respectively $(\bm{a}_{d})_{i}\equiv\langle s(t-d)x_{i}(t)\rangle$ and $C_{ij}\equiv\langle x_{i}(t)x_{j}(t)\rangle$ ($i,j$ indicate the indices of the readout neurons). Under the scaling assumption of model hyperparameters, Eq.(3), the inverse of the covariance matrix $C$ can be expanded using Neumann series expansion:
$C^{-1}=\sum_{n=0}^{\infty}D^{-1}\left(-\tilde{C}D^{-1}\right)^{n},$ (48)
where $D_{ij}\equiv\delta_{ij}\langle x_{i}(t)^{2}\rangle$ and $\tilde{C}_{ij}\equiv(1-\delta_{ij})\langle x_{i}(t)x_{j}(t)\rangle$.
This series expansion is ensured to converge when the spectrum norm is less than one, $\|\tilde{C}D^{-1}\|<1$. However, calculating the exact value of the spectrum norm is a challenging task. Employing the inequality $\|\tilde{C}D^{-1}\|\leq\|\tilde{C}D^{-1}\|_{F}$, where $\|\cdot\|_{F}$ denotes a Frobenius norm, we can obtain the sufficient condition for the convergence, $\|\tilde{C}D^{-1}\|_{F}<1$. The value of $\|\tilde{C}D^{-1}\|_{F}$ is calculated as
$\|\tilde{C}D^{-1}\|_{F}$ $=\sum_{i,j(i\neq j)}^{\alpha\sqrt{N}}\frac{\langle x_{i}x_{j}\rangle^{2}}{\langle x_{i}^{2}\rangle^{2}}=\alpha^{2}N\left[\frac{\langle x_{i}x_{j}\rangle^{2}_{(i\neq j)}}{\langle x_{i}^{2}\rangle^{2}}\right]=\alpha^{2}N\frac{[\langle x_{i}x_{j}\rangle^{2}]_{(i\neq j)}}{[\langle x_{i}^{2}\rangle]^{2}}=\frac{\alpha^{2}}{[\langle x_{i}^{2}\rangle]^{2}}\frac{g^{4}\langle\phi(x)^{2}\rangle_{*}^{2}+\tilde{\sigma}_{s}^{4}}{1-g^{4}\langle\phi^{\prime}(x)\rangle_{*}^{4}},$ (49)
where the second equality employs the self-averaging property, and the final equality utilizes Eq.(41). In addition, the third equality follows from Eq.(28). Consequently, we obtain the sufficient condition,
$\alpha^{2}<\frac{[\langle x_{i}^{2}\rangle]^{2}}{\tilde{\sigma}_{s}^{4}+([\langle x_{i}^{2}\rangle]-\sigma_{n}^{2})^{2}}\left(1-(g\langle\phi^{\prime}(x)\rangle_{*})^{4}\right),$ (50)
where we used the equation, $g^{2}\langle\phi(x)^{2}\rangle_{*}=[\langle x_{i}^{2}\rangle]-\sigma_{n}^{2}$, derived from the dynamical mean-field equation, Eq(23).
Substituting the series expansion expression for $C^{-1}$ into Eq.(47), we obtain
$M_{d}=\frac{\sqrt{N}}{\tilde{\sigma}_{s}^{2}}\sum_{i}^{\alpha\sqrt{N}}\frac{\langle s_{t-d}x_{i,t}\rangle^{2}}{\langle x_{i}^{2}\rangle}-\frac{\sqrt{N}}{\tilde{\sigma}_{s}^{2}}\sum_{\begin{subarray}{c}i,j\\ (i\neq j)\end{subarray}}^{\alpha\sqrt{N}}\frac{\langle s_{t-d}x_{i,t}\rangle\langle x_{i}x_{j}\rangle\langle s_{t-d}x_{j,t}\rangle}{\langle x_{i}^{2}\rangle\langle x_{j}^{2}\rangle}+\frac{\sqrt{N}}{\tilde{\sigma}_{s}^{2}}\sum_{\begin{subarray}{c}i,j,k\\ (i\neq j,j\neq k)\end{subarray}}^{\alpha\sqrt{N}}\frac{\langle s_{t-d}x_{i,t}\rangle\langle x_{i}x_{j}\rangle\langle x_{j}x_{k}\rangle\langle s_{t-d}x_{k,t}\rangle}{\langle x_{i}^{2}\rangle\langle x_{j}^{2}\rangle\langle x_{k}^{2}\rangle}-\cdots,$
where for simplicity we used the shorthand notation $x_{i,t}\equiv x_{i}(t)$, $s_{t}\equiv s(t)$, and $\langle x_{i}x_{j}\rangle\equiv\langle x_{i}(t)x_{j}(t)\rangle$.
In the large network size limit, we can assume self-averaging for $M_{d}$, that is, $\lim_{N\to\infty}M_{d}=[M_{d}]$, where the square bracket denotes the average over network realizations, known as ”quenched average” [19]. In addition, as neurons exhibit statistically identical dynamics, the quenched averaging operation eliminates the dependence of each term in Eq.(2) on neuron indices $i,j,k,\cdots$. Consequently, Eq.(B.1) can be reduced to
$[M_{d}]=\frac{\alpha N}{\tilde{\sigma}_{s}^{2}}\left[\frac{\langle s_{t-d}x_{i,t}\rangle^{2}}{\langle x_{i}^{2}\rangle}\right]-\frac{\alpha^{2}N^{\frac{3}{2}}}{\tilde{\sigma}_{s}^{2}}\left[\frac{\langle s_{t-d}x_{i,t}\rangle\langle x_{i}x_{j}\rangle\langle s_{t-d}x_{j,t}\rangle}{\langle x_{i}^{2}\rangle\langle x_{j}^{2}\rangle}\right]_{i\neq j}+\frac{\alpha^{3}N^{2}}{\tilde{\sigma}_{s}^{2}}\left[\frac{\langle s_{t-d}x_{i,t}\rangle\langle x_{i}x_{j}\rangle\langle x_{j}x_{k}\rangle\langle s_{t-d}x_{k,t}\rangle}{\langle x_{i}^{2}\rangle\langle x_{j}^{2}\rangle\langle x_{k}^{2}\rangle}\right]_{\begin{subarray}{c}i\neq j\\ j\neq k\end{subarray}}-\cdots.$
For the solution of memory capacity, it is enough to calculate each quenched averaging term in Eq.(B.1). We perform this calculation by the dynamical cavity method, similar to the approach described in Appendix.A. As noted there, for a reservoir RNN with discrete time dynamics, perturbations induced by the auxiliary neurons do not contribute to the leading order term, which simplifies the calculation.
To begin with, we calculate the first term. Employing Eq.(28) and introducing an auxiliary neuron indexed by 00, it is enough to calculate $[\langle s_{t-d}x_{0,t}\rangle^{2}]$ as:
$[\langle s_{t-d}x_{0,t}\rangle^{2}]$ $=\left[\left\langle s_{t-d}\left(\sum_{i}^{N}J_{0i}\phi_{i,t-1}+u_{0}s_{t}+\xi_{0,t}\right)\right\rangle^{2}\right]=\sum_{i,j}^{N}\left[J_{0i}J_{0j}\langle s_{t-d}\phi_{i,t-1}\rangle^{2}\right]+[u_{0}^{2}]\langle s_{t-d}s_{t}\rangle^{2}$ (53)
$=\sum_{i,j}^{N}\left[J_{0i}J_{0j}\right]\left[\langle s_{t-d}\phi_{i,t-1}\rangle^{2}\right]+[u_{0}^{2}]\langle s_{t-d}s_{t}\rangle^{2}=g^{2}[\langle s_{t-(d-1)}\phi_{i,t}\rangle^{2}]+\delta_{d,0}\frac{\tilde{\sigma}_{s}^{4}}{N},$
where we used the independence between $J_{0,i},J_{0,j}$ and $\phi_{i,t}$ in the third equation. Since $(s_{t},x_{i,t})$ are Gaussian variables, we apply Price’s theorem, resulting in
$[\langle s_{t-d}x_{0,t}\rangle^{2}]=g^{2}[\langle\phi^{\prime}_{i}\rangle^{2}][\langle s_{t-(d-1)}x_{0,t}\rangle^{2}]+\delta_{d,0}\frac{\tilde{\sigma}_{s}^{4}}{N},$
where we restored $[\langle s_{t-(d-1)}x_{i,t}\rangle^{2}]$ to $[\langle s_{t-(d-1)}x_{0,t}\rangle^{2}]$. This equation is a recurrence formula for $[\langle s_{t-d}x_{0,t}\rangle^{2}]$, whose solution is
$[\langle s_{t-d}x_{0,t}\rangle^{2}]=\frac{1}{N}\tilde{\sigma}_{s}^{4}\left(g\langle\phi^{\prime}(x)\rangle_{*}\right)^{2d}.$ (55)
Consequently, we obtain
[…]
Next, we move on to the calculation of the second term. We introduce the two auxiliary neurons indexed by 00 and $0^{\prime}$, obtaining for $d\geq 1$
$\left[\langle s_{t-d}x_{0,t}\rangle\langle x_{0}x_{0^{\prime}}\rangle\langle s_{t-d}x_{0^{\prime},t}\rangle\right]$ $=\left[\sum_{i}^{N}J_{0i}\langle s_{t-d}\phi_{i,t-1}\rangle\sum_{j,k}^{N}J_{0j}J_{0^{\prime}k}\langle\phi_{j}\phi_{k}\rangle\sum_{l}^{N}J_{0^{\prime}l}\langle\phi_{l,t-1}s_{t-d}\rangle\right]$
$=\sum_{i,j,k,l}^{N}\left[J_{0i}J_{0j}J_{0^{\prime}k}J_{0^{\prime}l}\langle s_{t-d}\phi_{i,t-1}\rangle\langle\phi_{j}\phi_{k}\rangle\langle\phi_{l,t-1}s_{t-d}\rangle\right]$
$=\sum_{i,j,k,l}^{N}\left[J_{0i}J_{0j}J_{0^{\prime}k}J_{0^{\prime}l}\right]\left[\langle s_{t-d}\phi_{i,t-1}\rangle\langle\phi_{j}\phi_{k}\rangle\langle\phi_{l,t-1}s_{t-d}\rangle\right]$
$=\frac{g^{4}}{N^{2}}\sum_{i,j}^{N}\left[\langle s_{t-d}\phi_{i,t-1}\rangle\langle\phi_{i}\phi_{j}\rangle\langle\phi_{j,t-1}s_{t-d}\rangle\right]$
$=\frac{g^{4}}{N}[\langle s_{t-(d-1)}\phi_{i,t}\rangle^{2}][\langle\phi_{i}^{2}\rangle]+g^{4}\left[\langle s_{t-(d-1)}\phi_{i,t}\rangle\langle\phi_{i}\phi_{j}\rangle\langle\phi_{j,t}s_{t-(d-1)}\rangle\right]_{(i\neq j)}$
$=\frac{g^{4}}{N}[\langle s_{t-(d-1)}\phi_{0,t}\rangle^{2}][\langle\phi_{i}^{2}\rangle]+g^{4}[\langle\phi^{\prime}(x_{i})\rangle^{2}]^{2}\left[\langle s_{t-(d-1)}x_{0,t}\rangle\langle x_{0}x_{0^{\prime}}\rangle\langle x_{0^{\prime},t}s_{t-(d-1)}\rangle\right],$
where the last equality follows from Price’s theorem. This equation is a recurrence formula for $\left[\langle s_{t-d}x_{0,t}\rangle\langle x_{0}x_{0^{\prime}}\rangle\langle s_{t-d}x_{0^{\prime},t}\rangle\right]$ with the initial condition ($d=0$),
$\left[\langle s_{t}x_{0,t}\rangle\langle x_{0}x_{0^{\prime}}\rangle\langle s_{t}x_{0^{\prime},t}\rangle\right]=\frac{\tilde{\sigma}_{s}^{4}}{N}[\langle x_{0}x_{0^{\prime}}\rangle u_{0}u_{0^{\prime}}]=\frac{\tilde{\sigma}_{s}^{4}}{N}\left[u_{0}u_{0^{\prime}}\left(\sum_{i,j}^{N}J_{0i}J_{0^{\prime}j}\langle\phi_{i}\phi_{j}\rangle+u_{0}u_{0^{\prime}}\langle s_{t}^{2}\rangle\right)\right]=\tilde{\sigma}_{s}^{6}N^{-3/2}.$ (58)
Thus, the second term of Eq.(B.1) for $d=1$ is $O(N^{-3/2})$. On the other hand, from Eq.(55), the first term of Eq.(B.1) for $d=1$ scales as $O(N^{-2})$. As a result, this term becomes subleading and negligible. Consequently, solving the recurrence formula yields
$\left[\langle s_{t-d}x_{0,t}\rangle\langle x_{0}x_{0^{\prime}}\rangle\langle s_{t-d}x_{0^{\prime},t}\rangle\right]=\frac{\tilde{\sigma}_{s}^{6}\left(g\langle\phi^{\prime}(x)\rangle_{*}\right)^{4d}}{N^{3/2}}+{\rm subleading\ terms}$ (59)
and thus
[…]
Subsequently, we proceed to calculate the third term in Eq.(B.1). It is enough to consider the case for $i\neq k$ because the terms for $i=k$ have the same order as those for $i\neq k$, but the number of $i=k$ terms, $O(N^{2})$, is much smaller than those of $i\neq k$ terms, $O(N^{3})$, making their overall contribution subleading. We introduce three auxiliary neurons, labeled 00, $0^{\prime}$, and $0^{\prime\prime}$, obtaining for $d\geq 1$
$\left[\langle s_{t-d}x_{0,t}\rangle\langle x_{0}x_{0^{\prime}}\rangle\langle x_{0^{\prime}}x_{0^{\prime\prime}}\rangle\langle x_{0^{\prime\prime},t}s_{t-d}\rangle\right]$ $=\left[\sum_{i}^{N}J_{0i}\langle s_{t-d}\phi_{i,t-1}\rangle\sum_{j,k}^{N}J_{0j}J_{0^{\prime}k}\langle\phi_{j}\phi_{k}\rangle\sum_{l,m}^{N}J_{0^{\prime}l}J_{0^{\prime\prime}m}\langle\phi_{l}\phi_{m}\rangle\sum_{n}^{N}J_{0^{\prime\prime}n}\langle\phi_{n,t-1}s_{t-d}\rangle\right]$ (61)
$=\frac{g^{6}}{N^{3}}\sum_{i,j,k}^{N}\left[\langle s_{t-d}\phi_{i,t-1}\rangle\langle\phi_{i}\phi_{j}\rangle\langle\phi_{j}\phi_{k}\rangle\langle\phi_{k,t-1}s_{t-d}\rangle\right]$
$=\frac{g^{6}}{N^{2}}[\langle\phi_{i}^{2}\rangle]^{2}[\langle s_{t-(d-1)}\phi_{i,t}\rangle^{2}]+\frac{2g^{6}}{N}[\langle\phi_{i}^{2}\rangle]\left[\langle s_{t-(d-1)}\phi_{0,t}\rangle\langle\phi_{0}\phi_{0^{\prime}}\rangle\langle\phi_{0^{\prime},t}s_{t-(d-1)}\rangle\right]$
$\hskip 60.0pt+g^{6}[\langle\phi^{\prime}_{i}\rangle^{2}]^{3}[\langle s_{t-(d-1)}x_{0,t}\rangle\langle\phi_{0}\phi_{0^{\prime}}\rangle\langle x_{0^{\prime}}x_{0^{\prime\prime}}\rangle\langle x_{0^{\prime\prime},t}s_{t-(d-1)}\rangle],$
where we used Price’s theorem in the last equality. This equation is a recurrence formula for $\left[\langle s_{t-d}x_{0,t}\rangle\langle x_{0}x_{0^{\prime}}\rangle\langle x_{0^{\prime}}x_{0^{\prime\prime}}\rangle\langle x_{0^{\prime\prime},t}s_{t-d}\rangle\right]$ with initial condition ($d=0$),
$\left[\langle s_{t}x_{0,t}\rangle\langle x_{0}x_{0^{\prime}}\rangle\langle x_{0^{\prime}}x_{0^{\prime\prime}}\rangle\langle x_{0^{\prime\prime},t}s_{t}\rangle\right]$ $=\frac{\tilde{\sigma}^{4}}{N}[\langle x_{0}x_{0^{\prime}}\rangle\langle x_{0^{\prime}}x_{0^{\prime\prime}}\rangle u_{0}u_{0^{\prime\prime}}]$ (62)
$=\frac{\tilde{\sigma}^{4}}{N}\left[u_{0}u_{0^{\prime\prime}}\left(\sum_{i,j}^{N}J_{0i}J_{0^{\prime}j}\langle\phi_{i}\phi_{j}\rangle+u_{0}u_{0^{\prime}}\langle(s_{t})^{2}\rangle\right)\left(\sum_{i,j}^{N}J_{0^{\prime}i}J_{0^{\prime\prime}j}\langle\phi_{i}\phi_{j}\rangle+u_{0^{\prime}}u_{0^{\prime\prime}}\langle(s_{t})^{2}\rangle\right)\right]$
$=\tilde{\sigma}^{8}N^{-2}.$
Therefore, the third term in Eq.(61) for $d=1$ is $O(1/N^{2})$, while the first and second term scale $O(1/N^{3})$ and $O(1/N^{5/2})$, respectively. As a result, only third term contributes to the leading order. Solving the recurrence formula, Eq.(61), to leading order, we obtain
$\left[\langle s_{t}x_{0,t}\rangle\langle x_{0}x_{0^{\prime}}\rangle\langle x_{0^{\prime}}x_{0^{\prime\prime}}\rangle\langle x_{0^{\prime\prime},t}s_{t}\rangle\right]=\frac{\tilde{\sigma}_{s}^{8}\left(g\langle\phi^{\prime}(x)\rangle_{*}\right)^{6d}}{N^{2}}+{\rm subleading\ terms}$ (63)
and thus
[…]
The analogous calculation can be applied to the remaining terms, and we finally obtain the analytical solution for $M_{d}$ and memory capacity:
$M_{d}=\sum_{n=0}^{\infty}(-1)^{n}\left\{\frac{\alpha\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}\left(g\langle\phi^{\prime}(x)\rangle_{*}\right)^{2d}\right\}^{n+1},\hskip 20.0ptMC=\sum_{d=0}^{\infty}\sum_{n=0}^{\infty}(-1)^{n}\left\{\frac{\alpha\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}\left(g\langle\phi^{\prime}(x)\rangle_{*}\right)^{2d}\right\}^{n+1}.$
B.2 Analytical solution of the decay rate, $\lim_{N\to\infty}r(L)$
Noting that by applying Eq.(B.1), the denominator in the definition of the decay rate, Eq.(11), can be calculated as
[…]
$=\frac{\alpha\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]\left(1-g^{2}\langle\phi^{\prime}(x)\rangle_{*}^{2}\right)},$
the decay rate of memory capacity can be analytically derived as
$\lim_{N\to\infty}r(L=\alpha\sqrt{N})=\frac{[\langle x_{i}^{2}\rangle]\left(1-g^{2}\langle\phi^{\prime}(x)\rangle_{*}^{2}\right)}{\alpha\tilde{\sigma}_{s}^{2}}\sum_{d=0}^{\infty}\sum_{n=0}^{\infty}(-1)^{n}\left\{\frac{\alpha\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}\left(g\langle\phi^{\prime}(x)\rangle_{*}\right)^{2d}\right\}^{n+1}.$ (67)
The two infinite summations, $\sum_{d=0}^{\infty}$ and $\sum_{n=1}^{\infty}$, can be swapped under the condition, Eq.(50), because the infinite series is absolutely convergent:
$\sum_{d=0}^{\infty}\sum_{n=0}^{\infty}\left\{\frac{\alpha\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}\left(g\langle\phi^{\prime}(x)\rangle_{*}\right)^{2d}\right\}^{n+1}=\sum_{d=0}^{\infty}\frac{\frac{\alpha\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}\left(g\langle\phi^{\prime}(x)\rangle_{*}\right)^{2d}}{1-\frac{\alpha\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}\left(g\langle\phi^{\prime}(x)\rangle_{*}\right)^{2d}}<\sum_{d=0}^{\infty}\frac{\alpha\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}\left(g\langle\phi^{\prime}(x)\rangle_{*}\right)^{2d}<\frac{\alpha\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}\frac{1}{1-(g\langle\phi^{\prime}(x)\rangle_{*})^{2}}<\infty.$
Here, in the first inequality, we apply the result $10<g\langle\phi^{\prime}(x)\rangle_{*}<$ 1 proven in Appendix.A.3, and $\frac{\alpha\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}<1$ as shown below:
$\frac{\alpha\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}<\frac{\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}\sqrt{\frac{[\langle x_{i}^{2}\rangle]^{2}}{\tilde{\sigma}_{s}^{4}+([\langle x_{i}^{2}\rangle]-\sigma_{n}^{2})^{2}}\left(1-(g\langle\phi^{\prime}(x)\rangle_{*})^{4}\right)}<\frac{\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}\sqrt{\frac{[\langle x_{i}^{2}\rangle]^{2}}{\tilde{\sigma}_{s}^{4}}}=1,$ (69)
where we used Eq.(50) in the first inequality.
Consequently, by interchanging the two infinite summations, the decay rate can be represented as
$\lim_{N\to\infty}r(L=\alpha\sqrt{N})=1-\sum_{n=1}^{\infty}(-1)^{n-1}\left(\frac{\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}\alpha\right)^{n}\frac{1-(g\langle\phi^{\prime}(x)\rangle_{*})^{2}}{1-(g\langle\phi^{\prime}(x)\rangle_{*})^{2n+2}},$ (70)
where the second term arises due to the neuronal correlations.
B.3 Decay rate for a linear RNN
We mention the decay rate of memory capacity for the reservoir RNN with a linear activation function, $\phi(x)=x$. In this case, the decay rate can be derived as
$\lim_{N\to\infty}r(L=\alpha\sqrt{N})=1-\sum_{n=1}^{\infty}(-1)^{n-1}\left(\frac{\tilde{\sigma}_{s}^{2}}{\sigma_{n}^{2}}\alpha\right)^{n}\frac{(1-g^{2})^{n+1}}{1-g^{2n+2}},$ (71)
where the value of $g$ must be less than one to ensure the stability of the reservoir RNN. Analogous to non-linear RNNs, the memory capacity of the linear RNN with neuronal noise exhibits sublinear scaling with respect to $L$, as illustrated in Fig.7. In contrast, for a linear RNN in the absence of noise, the memory capacity is known to be exactly $L$ [6], resulting in $r(L)=1$. It is intriguing that the presence or absence of noise greatly affects the scaling behavior of memory capacity with respect to $L$.
Here, we mention several points. First, the sublinear scaling of memory capacity for a linear RNN subject to neuronal noise is not accompanied by an increase in nonlinear computational capabilities, unlike in the case of a nonlinear RNN (Fig.5(a)(b)). This is simply because a linear RNN lacks the ability to perform nonlinear computations. Consequently, in a linear RNN, the sublinear scaling of memory capacity solely reflects a decline in overall computational performance caused by neuronal noise. Second, our theoretical analysis based on the Neumann series expansion is not applicable to the noise-free linear RNN. In this setting, the diagonal entries of covariance matrix $C$ are comparable to its non-diagonal entries, both scaling as $O(1/\sqrt{N})$, which prevents us from employing the Neumann series expansion.
[…]
Appendix C INFORMATION PROCESSING CAPACITY THEORY
C.1 Overview of IPC theory
We provide a concise overview of the IPC, including its definition and key properties. We consider a reservoir receiving input signals $\{s(t)\}_{t}$. The general task is formulated as optimizing output weights such that the reservoir’s output $z(t)$ approximates a given function of the inputs, $f[\cdots,s(t-1),s(t)]$. For input signals drawn i.i.d from a standard Gaussian distribution, the orthonormal basis functions spanning the Hilbert space containing $f$ are expressed as a formally infinite product of normalized Hermite polynomials:
$y_{\bm{d}}=\prod_{i\geq 0}\mathcal{H}_{d_{i}}\left(s(t-i)\right),$ (72)
where $\mathcal{H}_{d_{i}}(\cdot)$ ($d_{i}\geq 0$) denotes the normalized Hermite polynomial of degree $d_{i}$, and vector-form index $\bm{d}\equiv(d_{i})_{i\geq 0}$ has only a finite number of non-zero elements. Since $\mathcal{H}_{0}=1$, the product in Eq.(72) is effectively finite. The summation of $d_{i}$ is equivalent to the degree of the polynomial $y_{\bm{d}}$, defined as $deg(y_{\bm{d}})\equiv\sum_{i\geq 0}d_{i}$.
The Capacity for the reservoir to reconstruct the function of inputs, $f$, is defined as
$C_{T}[f]\equiv 1-\frac{{\rm{min}}_{\bm{w}}\langle(z(t)-f(t))^{2}\rangle_{T}}{\langle f(t)^{2}\rangle_{T}},$ (73)
where $\langle\cdot\rangle_{T}$ denotes the time average over simulation time $T$. It is proven that $0\leq C_{T}[f]\leq 1$ holds true for any function $f$. The total IPC is subsequently defined as the summation of capacities across all basis polynomials:
$IPC_{\rm{total}}\equiv\sum_{\bm{d}}C_{T}[y_{\bm{d}}].$ (74)
It has been established that $\lim_{T\to\infty}IPC_{\rm{total}}$ is equivalent to the number of readout units, $L$, provided that the reservoir satisfies echo state property [2], whereby the reservoir’s state is uniquely determined solely by input signals, regardless of its initial state.
The total IPC can be decomposed based on the degree of the basis polynomials. We define the IPC for degree $D$ as
$IPC_{D}\equiv\sum_{\begin{subarray}{c}\bm{d}\\ {\rm{s.t.}}\ deg(y_{\bm{d}})=D\end{subarray}}C_{T}[y_{\bm{d}}].$ (75)
Crucially, $IPC_{1}$ is identical to memory capacity, as $C_{T}[y_{\bm{d}}]$ for $deg(y_{\bm{d}})=1$ corresponds to Eq.(2). In contrast, $IPC_{D}$ for $D\geq 2$ represents a non-linear computational ability of the reservoir.
When calculating the capacities by numerical simulation, we must take care not to overestimate them, because for finite simulation time $T$, they are subject to a systematic positive error. Following the approach of Dambre et al. [7], we determine the threshold of the capacities, $\epsilon$, as
$\epsilon\equiv\frac{2\theta}{T},\ \ \theta\equiv\arg_{\theta}\left\{\mathbb{P}[\chi^{2}(L)\geq\theta]=p\right\},$ (76)
where $\chi^{2}(L)$ denotes a random variable that follows a Chi-squared distribution with $L$ degrees of freedom. The value of $p$ represents the probability that a truly zero capacity is incorrectly assessed as non-zero. We assume $C_{T}=0$ if $C_{T}$ is lower than a threshold value $\epsilon$. Following Dambre et al., we set $p=10^{-4}$ for all our numerical simulations of memory capacity ($=IPC_{1}$) and IPC, although the choice of $p$ has a negligible impact on our results.
Note that calculating the IPC values of our reservoir RNN model requires a minor correction to the time evolution equation, Eq.1, as
$x_{i}(t)=\sum_{j=1}^{N}J_{ij}\phi(x_{i}(t-1))+u_{i}\sigma_{s}s(t)+\xi_{i}(t),\ \ s(t)\sim\mathcal{N}(0,1),$ (77)
where the standard deviation of the input signals, $\sigma_{s}$, is now incorporated as a scaling factor of the standard Gaussian noise inputs. This correction is necessary because IPC is defined for a reservoir receiving standard Gaussian noise inputs. Importantly, this modification does not alter the statistical behavior or computational ability of the original RNN model.
C.2 Simulation results for IPC for an RNN with the ReLU activation function
Our theoretical framework for the analytical derivation of memory capacity cannot be applied to the RNNs with the ReLU activation function. This limitation arises because Price’s theorem (Eq.(37)) cannot be used for the ReLU function, as it is non-smooth at the origin. Therefore, we numerically obtain memory capacity and IPC of a ReLU RNN, as shown in Fig.8. Similarly to the RNNs employing the sigmoid function, the memory capacity ($IPC_{1}$) increases sublinearly with $L$, while non-linear computational capacities ($IPC_{D}$ for $D\geq 2$) emerge supralinearly and sequentially.
Note that unlike the sigmoid RNNs, the ReLU RNNs do not exhibit chaotic behavior. Instead, they diverge when the parameter $g$ exceeds a critical value, $g_{c}$. To demonstrate this, substituting $\phi(x)={\rm max}(0,x)$ into the dynamical mean-field equation (Appendix.A.1), we obtain
$[x_{i}(t+1)^{2}]=\frac{g^{2}}{2}[x_{i}(t)^{2}]+s(t+1)^{2}+\sigma_{n}^{2}+O(1/\sqrt{N}).$ (78)
This self-consistent equation admits a stable solution only when $g<g_{c}=\sqrt{2}$. For $g>g_{c}$, the system diverges regardless of the parameters $\sigma_{s}$ and $\sigma_{n}$. Consequently, the numerical simulations presented in Fig.8 are conducted with $g$ values strictly less than $\sqrt{2}$.
Figure 8: IPC values for varying the number of readout neurons $L$ of a reservoir RNN with the ReLU activation function $\phi(x)={\rm max}(0,x)$ and network size $N=1000$. The figure’s structure and content mirror Fig.5. The values of IPC for up to degree 5 are calculated. Simulation time length is $T=10^{4}$. The model parameters $(g,\sigma_{s},\sigma_{n})$ are $(1.0,1.0,0.0)$ for (a) and $(1.0,1.0,0.3)$ for (b). The RNN for (a) satisfies the echo state property since its positive maximum conditional Lyapunov exponent is negtive ($\lambda_{\rm max}=-0.35\pm 0.007$), while the property breaks down for (b) as the RNN receives neuronal noise.
C.3 Linear scaling of $IPC_{D}$ with $L$ for vanishing neuronal correlations
In the main text, we claim that each $IPC_{D}$ for $D\geq 1$ exhibits linear scaling with L if we ignore neuronal correlations. The following provides a concise proof of this claim. First, similarly to Eq.(4), the capacity for a basis function $y_{\bm{d}}$ can be expressed as
$C_{T}[y_{\bm{d}}]=\frac{\bm{a}^{\top}C^{-1}\bm{a}}{\langle y_{\bm{d}}(t)^{2}\rangle},$ (79)
where the elements of $\bm{a}\in\mathbb{R}^{L}$ are $a_{i}=\langle y_{\bm{d}}(t)x_{i}(t)\rangle$ and the matrix $C\in\mathbb{R}^{L\times L}$ is the covariance matrix of the readout neurons. Assuming neuronal correlations vanish, the capacity for any $y_{\bm{d}}$ is proportional to $L$ since
$C_{T}[y_{\bm{d}}]$ $\approx\left[C_{T}[y_{\bm{d}}]\right]\approx\left[\sum_{i}^{L}\frac{\langle y_{\bm{d}}(t)x_{i}(t)^{2}\rangle}{\langle x_{i}(t)^{2}\rangle\langle y_{\bm{d}}(t)^{2}\rangle}\right]=L\left[\frac{\langle y_{\bm{d}}(t)x_{i}(t)\rangle^{2}}{\langle x_{i}(t)^{2}\rangle\langle y_{\bm{d}}(t)^{2}\rangle}\right]\propto L,$ (80)
where we assume the self-averaging of the capacity in the first approximation equation, and in the second approximation equation, we ignore the correlation $\langle x_{i}x_{j}\rangle$.